OpenStudy (paki):

How we will find the square root of any large term like (97697.94) without calculator....

OpenStudy (anonymous):



OpenStudy (triciaal):

one method put the number in scientific notation then adjust to use even exponents of 10 the resulting numbers should be less than 100 and you can approximate the sq rts of those then expand with the corresponding factor of 10 example 97697.94 = 9.769794 *10^ 4 sqrt 97697.94 = sqrt 9.769794*sqrt 10^4 = 3 *10^2 = 300 depending on how accurate the results are needed incorporate the method above.. the method I was taught is different. can't remember right now but involved knowing the perfect squares and how close the desired number was to the perfect square number.